A firm uses 50 percent common stock, 40 percent debt, and 10 percent preferred stock. the cost of equity is 14 percent, the cost of preferred is 12 percent, and the pre-tax cost of debt is 7 percent. What is the WACC if the tax rate is 35 percent?
Answer #1

After-tax cost of debt=7*(1-tax rate)

=7*(1-0.35)=4.55%

WACC=Respective costs*Respective weight

=(0.5*14)+(0.4*4.55)+(0.1*12)

which is equal to

=10.02%